# Complex Domain Approach for Reversible Data Hiding and Homomorphic Encryption: General Framework and Application to Dispersed Data

### v1

@@ Abstract
I Introduction
@@
The ubiquitous generation and pervasive use of digital data stand as defining characteristics of the modern era. From vast networks of Internet of Things (IoT) devices and industrial sensors to intricate smart city infrastructures and critical healthcare systems, data fuels decision-making, automates processes, and enables unprecedented levels of connectivity. This reliance on data, often collected from distributed and potentially vulnerable environments, is critically linked to its trustworthiness [23]. Ensuring the confidentiality, integrity, authenticity, and privacy of information has become a priority, as compromised data can lead to erroneous conclusions, system failures, and severe breaches of privacy or security [14].
@@
Within this complex digital landscape, the concept of Reversible Data Hiding (RDH) [21, 13] emerges as a specialized and increasingly popular approach to data security. RDH involves embedding secret information (a “watermark”) directly into a host data item in such a way that, after extraction of the watermark, the original host data can be perfectly and exactly restored to its original state. This unique property, combining data embedding with guaranteed reversibility, makes RDH appropriate in sensitive applications such as medical imaging, legal documentation, and military intelligence, where any permanent alteration of the original data is unacceptable. Unlike traditional data encryption, which primarily ensures confidentiality, RDH focuses on providing intrinsic integrity, authenticity, and provenance, within the data itself.
@@
In the context of individual, scalar, or time-series numerical data, such as those originating from Wireless Sensor Networks (WSNs) [20], IoT devices [19], or critical financial and medical monitoring systems [2], RDH offers unique advantages. Foremost among these is its ability to ensure robust data integrity and authenticity at the granular level of single readings or sequential observations. By embedding a watermark that can be extracted and verified, receivers (e.g., a Data Collector in a WSN) can confirm that the data originated from a legitimate source and has not been tampered with in transit. This inherent capability to facilitate data provenance and traceability is particularly valuable, allowing for forensic analysis of data origin and processing paths in distributed and often resource-constrained environments.
@@
Beyond these general benefits, RDH offers distinct advantages compared to other commonly employed security primitives in certain contexts. Unlike cryptographic hashing or digital signatures, which typically append a separate authentication tag or signature to the data, RDH embeds the integrity verification mechanism directly within the data itself. This can simplify data handling and ensure the integrity proof travels jointly with the content. While conventional encryption ensures confidentiality, it does not directly provide data integrity; and Physical Unclonable Functions (PUFs) [9] focus on device authentication rather than data integrity. RDH uniquely combines content-dependent integrity (through watermark verification) with the guarantee of perfect original data restoration, a feature not offered by one-way hash functions or irreversible authentication tags. This intrinsic embedding can be particularly beneficial in constrained environments or specialized data formats where appending additional metadata is undesirable or challenging.
@@
In a parallel development within cryptography, Homomorphic Encryption (HE) [1] has emerged as a transformative technology for data privacy [10]. HE schemes allow computations to be performed directly on encrypted data without ever needing to decrypt it [24]. The output of such a computation is an encrypted result that, when decrypted, is identical to the result of the same computation performed on the unencrypted data. This property enables the processing of sensitive information by untrusted third parties (such as cloud servers) while ensuring that the data’s confidentiality remains fully protected throughout its lifecycle.
@@
I-A Plan and Contributions of the Paper
@@
The rest of the paper is organized as follows. Section II provides the necessary background on RDH and HE. Section III details the proposed complex-domain reversible watermarking scheme for both individual and aggregated data. Section IV presents the extension of ElGamal and Paillier encryption to the complex domain. Section V details the H[i][i]dden-EG and H[i][i]dden-AggP protocols for joint watermarking and encryption/aggregation, respectively. A comprehensive security analysis is provided in Section VI. Finally, Section VII concludes the paper and outlines future research directions.
II Background
This section provides the foundational concepts and related work that are essential for understanding the proposed framework. The core of this work lies at the intersection of two distinct but complementary fields of research: RDH and HE. While RDH offers powerful mechanisms for embedding information within data for integrity and provenance, HE provides thea unique capability to perform computations on encrypted data, thereby ensuring confidentiality. The following sections detail the relevant background on these two key areas.
II-A Reversible Data Hiding
Reversible Data Hiding [21, 13] is a specialized field within information hiding that focuses on embedding secret information into host data in such a way that the original host data can be perfectly and exactly recovered after the hidden data (watermark) has been extracted. This property is essential in applications that do not allow the permanent alteration of the original content.
Historically, RDH techniques have been extensively developed and applied primarily to multimedia content, including digital images, audio, and video [21]. Methods like Least Significant Bit (LSB) modification (when combined with original LSB storage) [5], Difference Expansion (DE) [22], Histogram Shifting (HS) [15], and Prediction-Error Expansion (PEE) [16] use the inherent statistical redundancy and the human perceptual system’s tolerance for slight alterations in these rich data formats. These approaches can embed information in a perceptually imperceptible manner while ensuring perfect reversibility, meaning the host data is faithfully restored to its original state after hidden data extraction.
@@ II-A Reversible Data Hiding
For this class of individual or dispersed data, existing reversible embedding methods tend to fall into a few categories. One common approach is Modulo/Remainder-based embedding, where the secret information $w$ (an integer in range $0,\ldots,q-1$) is embedded into a host integer $d$ using an operation like $d^{\prime}=q\cdot d+w$ [11]. A common special case involves embedding a single bit ($w\in\{0,1\}$) by setting $q=2$, resulting in $d^{\prime}=2d+w$ [12]. Upon extraction, $w$ is simply $d^{\prime}\,(\operatorname{\mathrm{mod}}q)$ and $d$ is $\lfloor d^{\prime}/q\rfloor$. When $q$ is a power of 2 ($q=2^{B}$), these extraction operations can be efficiently performed using bitwise logic (e.g., $w=d^{\prime}\text{ \& }(q-1)$ and $d=d^{\prime}\gg B$). While perfectly reversible and simple, a significant characteristic of this method is that the host data’s magnitude is directly scaled or inflated by a factor of $q$, making the numerical modification immediately obvious and leading to a poor mixing between the original data and the watermark.
@@
Another relevant approach for sequential data is Difference Expansion (DE), or its more general form, Prediction-Error Expansion [20]. These methods exploit relationships between consecutive data samples by computing prediction errors (e.g., $e_{k}=d_{k}-d_{k-1}$) and then embedding information by expanding these errors.
@@
II-B Homomorphic Encryption
Homomorphic Encryption [1] is a unique and powerful class of cryptographic schemes that allows a user to perform computations on encrypted data without first having to decrypt it. The result of these operations is an encrypted output which, when decrypted, is the same as if the operations had been performed on the original, unencrypted data. This remarkable capability is useful for maintaining data privacy and confidentiality when computations are outsourced to untrusted third parties, such as cloud services or remote data aggregators.
HE schemes are typically classified based on the types and number of operations they support. Partially Homomorphic Encryption (PHE) schemes, such as [6], support an unlimited number of a single type of operation, either addition or multiplication. Somewhat Homomorphic Encryption (SHE) schemes, which include the pioneering work by Boneh-Goh-Nissim (BGN) in [3], can perform a limited number of both additions and multiplications. Finally, Fully Homomorphic Encryption (FHE) schemes, such as Ring Learning with Errors (RLWE) [4], support an unlimited number of both operations, allowing for arbitrary computations on encrypted data. In this paper, two classical PHE schemes, ElGamal and Paillier, are applied for their distinct homomorphic properties that enable the proposed protocols for both individual data encryption and privacy-preserving data aggregation.
II-B1 ElGamal Encryption
ElGamal encryption [8] is an asymmetric key encryption algorithm based on the Diffie-Hellman key exchange. Its security relies on the difficulty of the discrete logarithm problem (DLP) in the multiplicative group of integers modulo a prime number.
@@ II-B1 ElGamal Encryption
  • Encryption: To encrypt a message $m\in\mathbb{Z}_{p}^{*}$, a random integer $b$, such that $1\leq b\leq p-1$, is selected as the ephemeral key. The ciphertext is then computed as $y_{1}=g^{b}\,(\operatorname{\mathrm{mod}}p)$ and $y_{2}=m\cdot K^{k}\,(\operatorname{\mathrm{mod}}p)$, resulting in the pair $(y_{1},y_{2})$. The encryption function can be denoted as follows:
@@
II-B2 Paillier Encryption
Paillier encryption is a probabilistic asymmetric encryption algorithm [17]. Unlike ElGamal, which is homomorphic with respect to multiplication, Paillier encryption is additively homomorphic, meaning that sums of plaintexts can be computed from their ciphertexts.
@@ II-B2 Paillier Encryption
      Compute $M_{P}=(L(g^{L_{P}}\,(\operatorname{\mathrm{mod}}n^{2})))^{-1}\,(\operatorname{\mathrm{mod}}n)$, where the function $L(x)$ is defined as $L(x)=\frac{x-1}{n}$.
@@
    The encryption function can be denoted as $\mathrm{Enc}_{\mathsf{K_{pub}}}(m,r)=g^{m}\cdot r^{n}\,(\operatorname{\mathrm{mod}}n^{2})$.
@@
    $\mathrm{Dec}_{\mathsf{K_{priv}}}(c_{1}\cdot c_{2}\,(\operatorname{\mathrm{mod}}n^{2}))=m_{1}+m_{2}\,(\operatorname{\mathrm{mod}}n).$
@@
III Reversible Watermarking Approach on the Complex Domain
@@ III-A Basic Definitions
Definition 1.
Fine-grained original data: $d_{j,k}$ is the scalar data produced by the $j$-th device (e.g., a sensor) in the $k$-th time sample (time $t_{k}$).
@@
Definition 2.
Watermark: $w_{k}$ is the watermark to be embedded at time $k$ in the fine-grained data $d_{j,k}$. Similarly, we assume $w_{k}\in\mathbb{Z}$.
@@
The only assumption we are making with respect to the watermark $w_{k}$ is to be an integer. Hence, any limitations in a practical deployment (e.g., a given number of bits) would be caused by the implementation details, such as the number of bits used for representing integers in a particular programming language. $\triangle$
Definition 3.
Reversible embedding and blind extraction functions: The embedding process is carried out using an embedding function $E_{w}$ in the following way:
@@
Note that the extraction is blind, since the original data is not used in the function.
Definition 4.
Aggregated data and watermarked aggregated data: The aggregated data $S_{k}$, for the devices $j=1,2,\dots,N$ in the $k$-th time sample is defined as follows:
@@
Definition 5.
Reversible blind aggregated data and watermark extraction: An aggregated data and watermark extraction function $R^{\mathrm{Agg}}_{w}$ allows extracting both $S_{k}$ and $w_{k}$ from $S^{\prime}_{k}$ as follows:
@@
III-B Proposed Reversible Watermarking Scheme
@@
As detailed in Definitions 1 and 2, both the data and the watermark are assumed to be integers and, hence, the complex number $\delta_{i,j}$ is indeed a Gaussian integer. $\triangle$
@@
$\delta^{\prime}_{j,k}$ is, again, a Gaussian integer, since it is the product of two Gaussian integers ($\lambda_{k}$ and $\delta_{i,j}$). Hence, the real and the imaginary parts of the watermarked data are integers. This is particularly interesting if we want to carry out further operations on this data, such as public-key encryption. $\triangle$
@@
While this is true, the idea behind the proposed scheme is to exploit the algebraic operations that are already available for complex numbers. There are libraries for complex number operations in most programming languages and, hence, they can be directly applied in any practical implementation of the method. This prevents the explicit use of matrices and matrix inverses, which may be cumbersome. Obviously, the method can be further generalized in its vector form, as discussed in the Appendices (supplemental material). $\triangle$
@@
Expression 2 perfectly matches the formulation given for $R_{w}$ in Definition 3, since it takes, as input parameters, $\delta^{\prime}_{j,k}$ (analogous to $d^{\prime}_{j,k}$) and $K_{w}=\lambda_{k}$. $\triangle$
@@
The proposed method is vulnerable against Gaussian integer factorization. Even with just one eavesdropped value, a malicious observer may try to compute the Guassian integer factors of the watermarked data and obtain knowledge about the original data, the watermark or the watermarking key. Consequently, for practical applications, it is not recommended to rely solely on the challenge $\lambda_{k}$ as a security measure. The scheme can be combined with random numbers and/or cryptographic functions for increased security and privacy, as shown in the protocols proposed in Section V. $\triangle$
@@
Hence, an eavesdropper who accessed $\delta^{\prime}_{j,k}$ would not be able to determine the data or the watermark directly, but may try to infer information from Guassian integer factorization, as discussed in Remark 7. In addition, if the eavesdropper had access to the watermarked values corresponding to different time samples ($k$) or different devices ($j$), they could try to obtain even more information about $\lambda_{k}$, especially if it is not changed frequently. $\triangle$
III-C Data Aggregation

### v2

@@ Abstract
Keywords: Reversible Data Hiding, Complex Domain, Gaussian Integers, Homomorphic Encryption, Data Aggregation, Dispersed Data, Paillier, ElGamal, Data Integrity
1 Introduction
@@
The ubiquitous generation and pervasive use of digital data stand as defining characteristics of the modern era. From vast networks of Internet of Things (IoT) devices and industrial sensors to intricate smart city infrastructures and critical healthcare systems, data fuels decision-making, automates processes, and enables unprecedented levels of connectivity. This reliance on data, often collected from distributed and potentially vulnerable environments, is critically linked to its trustworthiness (Yan et al., 2014). Ensuring the confidentiality, integrity, authenticity, and privacy of information has become a priority, as compromised data can lead to erroneous conclusions, system failures, and severe breaches of privacy or security (Liang et al., 2016).
@@
Within this complex digital landscape, the concept of Reversible Data Hiding (RDH) (Shi et al., 2016; Kumar et al., 2022) emerges as a specialized and increasingly popular approach to data security. RDH involves embedding secret information (a “watermark”) directly into a host data item in such a way that, after extraction of the watermark, the original host data can be perfectly and exactly restored to its original state. This unique property, combining data embedding with guaranteed reversibility, makes RDH appropriate in sensitive applications such as medical imaging, legal documentation, and military intelligence, where any permanent alteration of the original data is unacceptable. Unlike traditional data encryption, which primarily ensures confidentiality, RDH focuses on providing intrinsic integrity, authenticity, and provenance, within the data itself.
@@
In the context of individual, scalar, or time-series numerical data, such as those originating from Wireless Sensor Networks (WSNs) (Shi and Xiao, 2013), IoT devices (Rahman et al., 2021), or critical financial and medical monitoring systems (Bhalerao et al., 2019), RDH offers unique advantages. Foremost among these is its ability to ensure robust data integrity and authenticity at the granular level of single readings or sequential observations. By embedding a watermark that can be extracted and verified, receivers (e.g., a Data Collector in a WSN) can confirm that the data originated from a legitimate source and has not been tampered with in transit. This inherent capability to facilitate data provenance and traceability is particularly valuable, allowing for forensic analysis of data origin and processing paths in distributed and often resource-constrained environments.
@@
Beyond these general benefits, RDH offers distinct advantages compared to other commonly employed security primitives in certain contexts. Unlike cryptographic hashing or digital signatures, which typically append a separate authentication tag or signature to the data, RDH embeds the integrity verification mechanism directly within the data itself. This can simplify data handling and ensure the integrity proof travels jointly with the content. While conventional encryption ensures confidentiality, it does not directly provide data integrity; and Physical Unclonable Functions (PUFs) (Gao et al., 2020) focus on device authentication rather than data integrity. RDH uniquely combines content-dependent integrity (through watermark verification) with the guarantee of perfect original data restoration, a feature not offered by one-way hash functions or irreversible authentication tags. This intrinsic embedding can be particularly beneficial in constrained environments or specialized data formats where appending additional metadata is undesirable or challenging.
@@
In a parallel development within cryptography, Homomorphic Encryption (HE) (Acar et al., 2018) has emerged as a transformative technology for data privacy (Iezzi, 2020). HE schemes allow computations to be performed directly on encrypted data without ever needing to decrypt it (Zhang et al., 2020). The output of such a computation is an encrypted result that, when decrypted, is identical to the result of the same computation performed on the unencrypted data. This property enables the processing of sensitive information by untrusted third parties (such as cloud servers) while ensuring that the data’s confidentiality remains fully protected throughout its lifecycle.
@@
1.1 Plan and Contributions of the Paper
@@
The rest of the paper is organized as follows. Section 2 provides the necessary background on RDH and HE. Section 3 details the proposed complex-domain reversible watermarking scheme for both individual and aggregated data. Section 4 presents the extension of ElGamal and Paillier encryption to the complex domain. Section 5 details the H[i][i]dden-EG and H[i][i]dden-AggP protocols for joint watermarking and encryption/aggregation, respectively. A comprehensive formal security and privacy analysis of the proposed protocols is provided in Section 6. Finally, Section 7 concludes the paper and outlines future research directions.
2 Background
This section provides the foundational concepts and related work that are essential for understanding the proposed framework. The core of this work lies at the intersection of two distinct but complementary fields of research: RDH and HE. While RDH offers powerful mechanisms for embedding information within data for integrity and provenance, HE provides the unique capability to perform computations on encrypted data, thereby ensuring confidentiality. The following sections detail the relevant background on these two key areas.
2.1 Reversible Data Hiding
Reversible Data Hiding (Shi et al., 2016; Kumar et al., 2022) is a specialized field within information hiding that focuses on embedding secret information into host data in such a way that the original host data can be perfectly and exactly recovered after the hidden data (watermark) has been extracted. This property is essential in applications that do not allow the permanent alteration of the original content.
Historically, RDH techniques have been extensively developed and applied primarily to multimedia content, including digital images, audio, and video (Shi et al., 2016). Methods like Least Significant Bit (LSB) modification (when combined with original LSB storage) (Celik et al., 2002), Difference Expansion (DE) (Tian, 2003), Histogram Shifting (HS) (Ni et al., 2006), and Prediction-Error Expansion (PEE) (Ou et al., 2013) use the inherent statistical redundancy and the human perceptual system’s tolerance for slight alterations in these rich data formats. These approaches can embed information in an imperceptible mmanner while ensuring perfect reversibility, meaning the host data is faithfully restored to its original state after hidden data extraction.
@@ 2.1 Reversible Data Hiding
For this class of individual or dispersed data, existing reversible embedding methods tend to fall into a few categories. One common approach is Modulo or Remainder-based embedding, where the secret information $w$ (an integer in range $0,\ldots,q-1$) is embedded into a host integer $d$ using an operation like $d^{\prime}=q\cdot d+w$ (Kabir et al., 2025). A common special case involves embedding a single bit ($w\in\{0,1\}$) by setting $q=2$, resulting in $d^{\prime}=2d+w$ (Kabir et al., 2024). Upon extraction, $w$ is simply $d^{\prime}\,(\operatorname{\mathrm{mod}}q)$ and $d$ is $\lfloor d^{\prime}/q\rfloor$. When $q$ is a power of 2 ($q=2^{B}$), these extraction operations can be efficiently performed using bitwise logic (e.g., $w=d^{\prime}\text{ \& }(q-1)$ and $d=d^{\prime}\gg B$).
While perfectly reversible and simple, a significant characteristic of this method is that the host data’s magnitude is directly scaled or inflated by a factor of $q$, making the numerical modification immediately obvious and leading to a poor mixing between the original data and the watermark.
@@
Another relevant approach for sequential data is Difference Expansion (DE), or its more general form, Prediction-Error Expansion (Shi and Xiao, 2013). These methods exploit relationships between consecutive data samples by computing prediction errors (e.g., $e_{k}=d_{k}-d_{k-1}$) and then embedding information by expanding these errors.
@@
2.2 Homomorphic Encryption
Homomorphic Encryption (Acar et al., 2018) is a unique and powerful class of cryptographic schemes that allows a user to perform computations on encrypted data without first having to decrypt it. The result of these operations is an encrypted output which, when decrypted, is the same as if the operations had been performed on the original, unencrypted data. This remarkable capability is useful for maintaining data privacy and confidentiality when computations are outsourced to untrusted third parties, such as cloud services or remote data aggregators.
HE schemes are typically classified based on the types and number of operations they support. Partially Homomorphic Encryption (PHE) schemes, such as (Cominetti and Simplicio, 2020), support an unlimited number of a single type of operation, either addition or multiplication. Somewhat Homomorphic Encryption (SHE) schemes, which include the pioneering work by Boneh-Goh-Nissim (BGN) in (Boneh et al., 2005), can perform a limited number of both additions and multiplications. Finally, Fully Homomorphic Encryption (FHE) schemes, such as Ring Learning with Errors (RLWE) (Brakerski and Vaikuntanathan, 2011), support an unlimited number of both operations, allowing for arbitrary computations on encrypted data. In this paper, two classical PHE schemes, ElGamal (1985) and Paillier (1999), are applied for their distinct homomorphic properties that enable the proposed protocols for both individual data encryption and privacy-preserving data aggregation.
2.2.1 ElGamal Encryption
ElGamal encryption (ElGamal, 1985) is an asymmetric key encryption algorithm based on the Diffie-Hellman key exchange scheme. Its security relies on the difficulty of the discrete logarithm problem (DLP) in the multiplicative group of integers modulo a prime number.
@@ 2.2.1 ElGamal Encryption
  • Encryption: To encrypt a message $m\in\mathbb{Z}_{p}^{*}$, a random integer $b$, such that $1\leq b\leq p-1$, is selected as the ephemeral key. The ciphertext is then computed as $y_{1}=g^{b}\,(\operatorname{\mathrm{mod}}p)$ and $y_{2}=m\cdot K^{b}\,(\operatorname{\mathrm{mod}}p)$, resulting in the pair $(y_{1},y_{2})$. The encryption function can be denoted as follows:
@@
2.2.2 Paillier Encryption
Paillier (1999) is a probabilistic asymmetric encryption algorithm. Unlike ElGamal, which is homomorphic with respect to multiplication, Paillier encryption is additively homomorphic, meaning that sums of plaintexts can be computed from their ciphertexts.
@@ 2.2.2 Paillier Encryption
      Compute $M_{P}=(L(g^{L_{P}}\,(\operatorname{\mathrm{mod}}n^{2})))^{-1}\,(\operatorname{\mathrm{mod}}n)$, where the function $L(x)$ is defined as
      $L(x)=\frac{x-1}{n}.$
@@
    The encryption function can be denoted as
    $\mathrm{Enc}_{\mathsf{K_{pub}}}(m,r)=g^{m}\cdot r^{n}\,(\operatorname{\mathrm{mod}}n^{2}).$
@@
    $\mathrm{Dec}_{\mathsf{K_{priv}}}(c_{1}\cdot c_{2}\,(\operatorname{\mathrm{mod}}n^{2}))=(m_{1}+m_{2})\,(\operatorname{\mathrm{mod}}n).$
@@
3 Reversible Watermarking Approach on the Complex Domain
@@ 3.1 Basic Definitions
Definition 1 (Fine-grained original data).
$d_{j,k}$ is the scalar data produced by the $j$-th device (e.g., a sensor) in the $k$-th time sample (time $t_{k}$).
@@
Definition 2 (Watermark).
$w_{k}$ is the watermark to be embedded at time $k$ in the fine-grained data $d_{j,k}$. Similarly, we assume $w_{k}\in\mathbb{Z}$.
@@
The only assumption we are making with respect to the watermark $w_{k}$ is to be an integer. Hence, any limitations in a practical application (e.g., a given number of bits) would be caused by the implementation details, such as the number of bits used for representing integers in a particular programming language. $\triangle$
Definition 3 (Reversible embedding and blind extraction functions).
The embedding process is carried out using an embedding function $E_{w}$ in the following way:
@@
Note that the extraction is assumed to be blind, since the original data is not used in the function.
Remark 2.
Unlike reversible data hiding for multimedia content or highly redundant data —where source redundancy allows for techniques such as difference expansion, histogram shifting, or prediction error expansion to hide information without necessarily increasing carrier size— embedding a watermark $w_{k}$ in a single scalar $d_{j,k}$ such that both can be perfectly recovered requires data expansion. For instance, as seen in Kabir et al. (2025), embedding a $B$-bit watermark using an operation like $d^{\prime}_{j,k}=2^{B}\cdot d_{j,k}+w_{k}$ increases the bit-length of the original data by exactly $B$ bits. From an information-theoretic perspective, if the host data consists of $A$ bits and the watermark comprises $B$ bits, and both are independent and occupy their full possible ranges, the combined state space consists of $2^{A+B}$ distinct values. In such cases, the result must be encoded with at least $A+B$ bits for perfect recovery. While “slack” in a data’s restricted range (e.g., values $d_{j,k}\in[101,200]$ encoded in a 1-byte container) can theoretically host bits without expansion, scalar reversible embedding of independent, full-range variables fundamentally requires increasing the bit-length of the transmitted values. $\triangle$
Definition 4 (Aggregated data and watermarked aggregated data).
The aggregated data $S_{k}$, for the devices $j=1,2,\dots,N$ in the $k$-th time sample is defined as follows:
@@
Definition 5 (Reversible blind aggregated data and watermark extraction).
An aggregated data and watermark extraction function $R^{\mathrm{Agg}}_{w}$ allows extracting both $S_{k}$ and $w_{k}$ from $S^{\prime}_{k}$ as follows:
@@
3.2 Proposed Reversible Watermarking Scheme
@@
As detailed in Definitions 1 and 2, both the data and the watermark are assumed to be integers and, hence, the complex number $\delta_{j,k}$ is indeed a Gaussian integer. $\triangle$
@@
$\delta^{\prime}_{j,k}$ is, again, a Gaussian integer, since it is the product of two Gaussian integers ($\lambda_{k}$ and $\delta_{j,k}$). Hence, the real and the imaginary parts of the watermarked data are integers. This is particularly interesting if we want to carry out further operations on this data, such as public-key encryption. $\triangle$
@@
While this is true, the idea behind the proposed scheme is to exploit the algebraic operations that are already available for complex numbers. There are libraries for complex number operations in most programming languages and, hence, they can be directly applied in any practical implementation of the method. This prevents the explicit use of matrices and matrix inverses, which may be cumbersome. Obviously, the method can be further generalized in its vector form, as discussed in Section 3.4. $\triangle$
@@
Expression (2) perfectly matches the formulation given for $R_{w}$ in Definition 3, since it takes, as input parameters, $\delta^{\prime}_{j,k}$ (analogous to $d^{\prime}_{j,k}$) and $K_{w}=\lambda_{k}$. $\triangle$
@@
The proposed method is vulnerable against Gaussian integer factorization. Even with just one eavesdropped value, a malicious observer may try to compute the Gaussian integer factors of the watermarked data and obtain knowledge about the original data, the watermark or the watermarking key. Consequently, for practical applications, it is not recommended to rely solely on the challenge $\lambda_{k}$ as a security measure. The scheme can be combined with random numbers and/or cryptographic functions for increased security and privacy, as shown in the protocols proposed in Section 5.
$\triangle$
@@
Hence, an eavesdropper who accessed $\delta^{\prime}_{j,k}$ would not be able to determine the data or the watermark directly, but may try to infer information from Gaussian integer factorization, as discussed in Remark 8. In addition, if the eavesdropper had access to the watermarked values corresponding to different time samples ($k$) or different devices ($j$), they could try to obtain even more information about $\lambda_{k}$, especially if it is not changed frequently.
$\triangle$
3.3 Data Aggregation